# JointMotion: Joint Self-supervision for Joint Motion Prediction

### v1

@@ 7 Conclusion
In this work, we introduced a SSL method for joint motion prediction of all agents within a traffic scene. Our SSL framework integrates scene-level and agent-level objectives that operate complementarily, enhancing the training speed and accuracy of motion prediction models. Notably, our approach outperforms recent contrastive and autoencoding methods previously utilized for pre-training in motion prediction. Moreover, our objectives enable effective transfer learning between the Waymo Open Motion and the Argoverse 2 Forecasting datasets. Our evaluations demonstrate significant performance improvements over non-ensembling state-of-the-art joint prediction methods or joint-prediction variants of marginal-prediction architectures, underscoring the robustness and effectiveness of our proposed method.

### v2

@@ 5 Conclusion
In this work, we introduced a SSL method for joint motion prediction of multiple agents within a traffic scene. Our SSL framework integrates scene-level and instance-level objectives that operate complementarily, enhancing the training speed and accuracy of motion prediction models. Notably, JointMotion outperforms recent contrastive and autoencoding methods for pre-training in motion prediction. Moreover, our method generalizes across different environment representations, information fusion mechanisms, and dataset characteristics. Our evaluations demonstrate significant performance improvements over non-ensembling joint prediction methods or joint-prediction variants of marginal-prediction architectures, underscoring the robustness and effectiveness of our proposed method.
Limitations. Our method is not as label-efficient as SSL methods in computer vision, which generate supervisory signals from raw images (e.g., [3]). Our pre-training does not include the downstream labels for motion prediction (i.e., future motion), but relies on past motion and map data.
Acknowledgments
The research leading to these results is funded by the German Federal Ministry for Economic Affairs and Climate Action within the project “NXT GEN AI METHODS – Generative Methoden für Perzeption, Prädiktion und Planung”. Furthermore, we acknowledge the financial support by the German Federal Ministry of Education and Research (BMBF) within the project HAIBrid (FKZ 01IS21096A).
Appendix
Evaluation metrics. Following [41], we use the mean average precision (mAP), the average displacement error (minADE), the final displacement error (minFDE), miss rate (MR), and overlap rate (OR) to evaluate motion predictions. All metrics are computed using the minimum mode for $k=6$ modes. Accordingly, the metrics for the mode closest to the ground truth are measured. We use the official challenge website to compute metrics on WOMD. For the AV2 dataset, we use the evaluation software provided by [21]. Therefore, joint modes (best scene-wide mode) are evaluated on the interactive splits and marginal modes (best mode for each agent individually) on the regular splits.
The Waymo Open Motion Dataset (WOMD) [41] is comprised of over 1.1 million data points extracted from 103,000 urban or suburban driving scenarios, spanning 20 seconds each. The state of object-agents includes attributes like position, dimensions, velocity, acceleration, orientation, and angular velocity. Each data point captures 1 second of past followed by 8 seconds of future data. We resample this time interval with 10Hz.